# Confidence on the Focal: Conformal Prediction with Selection-Conditional Coverage

### v1

@@ 4.1.1 Top-K selection
For this selection rule, the reference set of each $j\in\widehat{{\mathcal{S}}}_{\textnormal{topk}}$ takes a unified form:
^$\widehat{{\mathcal{R}}}_{\textnormal{topk}}\,:=\,\{i\in{\mathcal{I}}_{\textnormal{calib}}:S_{i}>T_{\textnormal{topk}}\}.$ (13)
Intuitively, $T_{\textnormal{topk}}$ refers to the $(m-K)$-th smallest score among the test units; when $S_{n+j}$ is among the top $K$, replacing $S_{n+j}$ with another score that is greater than $T_{\textnormal{topk}}$ does not change the selection outcome. This intuition is formalized in the following proposition, with the proof deferred to Appendix B.5.
@@
For the top-K selection rule, for any $j\in[m]$ and any selection taxonomy $\mathfrak{S}\in 2^{[m]}$, such that $j\in\widehat{{\mathcal{S}}}_{\textnormal{topk}}$ and $\widehat{{\mathcal{S}}}_{\textnormal{topk}}\in\mathfrak{S}$, the reference set obeys ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}_{\textnormal{topk}}$ for all $y\in{\mathcal{Y}}$.
Computing ^$\widehat{{\mathcal{R}}}_{\textnormal{topk}}$ as a universal reference set reduces the computation complexity of the reference construction step in Algorithm 1 to $O(\max(m,n))$. It is also worth mentioning that the definition of reference set here is agnostic to the choice of $\mathfrak{S}$.
4.1.2 Selection based on joint quantiles
In some cases, practitioners may want to select a test unit $n+j$ if its score $S_{n+j}$ greater than the $q$-th quantile of the calibration and test scores. Equivalently, we could write the selection set as
$\widehat{{\mathcal{S}}}_{\textnormal{cq}}=\{j\in[m]:S_{n+j}>T_{\textnormal{jq}}\},\quad\text{where}\quad T_{\textnormal{jq}}=\inf\Big{\{}t\in{\mathcal{T}}:\textstyle{\sum_{i=1}^{m+n}}\mathds{1}\{S_{i}\leq t\}\geq q(n+m)\Big{\}}$
By definition, the selection threshold $T_{\textnormal{jq}}$ is invariant to the permutation of the calibration and test scores, and it is straightforward to check that for all $j\in\widehat{{\mathcal{S}}}$, the reference set is in the form of
^$\widehat{{\mathcal{R}}}_{\textnormal{jq}}=\{i\in{\mathcal{I}}_{\textnormal{calib}}:S_{i}>T_{\textnormal{jq}}\}.$
Using ^$\widehat{{\mathcal{R}}}_{jq}$ reduces the computation complexity of the reference set in Algorithm 1 to $O(\max(m,n))$.
4.1.3 Selection based on calibration quantiles
Consider now a selection rule, where test unit $n+j$ is selected if its score $S_{n+j}$ is greater than the $q$-th quantile of of the calibration scores. Equivalently, we could write the selection set as
$\widehat{{\mathcal{S}}}_{\textnormal{cq}}=\big{\{}j\in[m]:S_{n+j}>T_{\textnormal{cq}}\big{\}},\text{ where }T_{\textnormal{cq}}=\inf\Big{\{}t\in{\mathcal{T}}:\textstyle{\sum_{i=1}^{n}}\mathds{1}\{S_{i}\leq t\}\geq qn\Big{\}}.$ (14)
For each $j\in\widehat{{\mathcal{S}}}$, the refernence set consists of all calibration units in the top $q$-th quantile, i.e.,
^$\widehat{{\mathcal{R}}}_{\textnormal{cq}}=\{i\in{\mathcal{I}}_{\textnormal{calib}}:S_{i}>T_{\textnormal{cq}}\}.$ (15)
The validity of ^$\widehat{{\mathcal{R}}}_{\textnormal{cq}}$ as the referenece set is established below, whose proof is deferred to appendix B.6.
@@ 4.1.3 Selection based on calibration quantiles
For the selection rule (14), for any $j\in[m]$ and any selection taxonomy $\mathfrak{S}$, such that $j\in\widehat{{\mathcal{S}}}$ and $\widehat{{\mathcal{S}}}\in\mathfrak{S}$, it holds that ^$\widehat{{\mathcal{R}}}_{\textnormal{cq}}=\widehat{{\mathcal{R}}}_{n+j}(y)$ defined in (7) for all $y\in{\mathcal{Y}}$.

### v3

@@ B.1 Top-K selection
In this case, the reference set for each $j\in\widehat{{\mathcal{S}}}_{\textnormal{topk}}$ takes a unified form:
^$\widehat{{\mathcal{R}}}_{\textnormal{topk}}\,:=\,\{i\in[n]:S_{i}>T_{\textnormal{topk}}\},$ (26)
which is also agnostic to the choice of $\mathfrak{S}$. Intuitively, $T_{\textnormal{topk}}$ refers to the $(m-K)$-th smallest score among the test units; when $S_{n+j}$ is among the top $K$, replacing $S_{n+j}$ with another score above $T_{\textnormal{topk}}$ does not change the selection outcome. This intuition is formalized in the following proposition, with its proof deferred to Supplementary Section D.7.
@@
With the top-K selection rule, for any $j\in[m]$ and any selection taxonomy $\mathfrak{S}\in 2^{[m]}$ such that $j\in\widehat{{\mathcal{S}}}_{\textnormal{topk}}$ and $\widehat{{\mathcal{S}}}_{\textnormal{topk}}\in\mathfrak{S}$, the reference set obeys ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}_{\textnormal{topk}}$ for all $y\in{\mathcal{Y}}$.
Computing ^$\widehat{{\mathcal{R}}}_{\textnormal{topk}}$ as a universal reference set reduces the computation complexity of the reference construction step in Algorithm 1 to $O(\max(m,n))$.
B.2 Selection based on joint quantiles
When selected units are those whose scores are above the $q$-th quantile of calibration and test scores, the selection set can be written as
$\widehat{{\mathcal{S}}}_{\textnormal{jq}}=\{j\in[m]:S_{n+j}>T_{\textnormal{jq}}\},\quad\text{where}\quad T_{\textnormal{jq}}=\inf\Big{\{}t\in{\mathcal{T}}:\textstyle{\sum_{i=1}^{m+n}}\mathds{1}\{S_{i}\leq t\}\geq q(n+m)\Big{\}}$
By definition, the selection threshold $T_{\textnormal{jq}}$ is invariant to the permutation of the calibration and test scores, and it is straightforward to check that for all $j\in\widehat{{\mathcal{S}}}_{\textnormal{jq}}$, the reference set is
^$\widehat{{\mathcal{R}}}_{\textnormal{jq}}=\{i\in[n]:S_{i}>T_{\textnormal{jq}}\}.$
Replacing ^$\widehat{{\mathcal{R}}}_{n+j}$’s with ^$\widehat{{\mathcal{R}}}_{jq}$ reduces the computation complexity of the reference set in Algorithm 1 to $O(\max(m,n))$.
B.2.1 Selection based on calibration quantiles
When selected units are those whose scores are greater than the $q$-th quantile of the calibration scores, the selection set is equivalently
$\widehat{{\mathcal{S}}}_{\textnormal{cq}}=\big{\{}j\in[m]:S_{n+j}>T_{\textnormal{cq}}\big{\}},\quad\text{ where}\quad T_{\textnormal{cq}}=\inf\Big{\{}t\in{\mathcal{T}}:\textstyle{\sum_{i=1}^{n}}\mathds{1}\{S_{i}\leq t\}\geq qn\Big{\}}.$ (27)
For each $j\in\widehat{{\mathcal{S}}}_{\textnormal{cq}}$, the reference set consists of all calibration units in the top $q$-th quantile, i.e.,
^$\widehat{{\mathcal{R}}}_{\textnormal{cq}}=\{i\in[n]:S_{i}>T_{\textnormal{cq}}\}.$ (28)
The validity of ^$\widehat{{\mathcal{R}}}_{\textnormal{cq}}$ as the reference set is established below, whose proof is deferred to Supplementary Section D.8.
@@ B.2.1 Selection based on calibration quantiles
For the selection rule specified in (27), for any $j\in[m]$ and any selection taxonomy $\mathfrak{S}$ such that $j\in\widehat{{\mathcal{S}}}$ and $\widehat{{\mathcal{S}}}\in\mathfrak{S}$, it holds that ^$\widehat{{\mathcal{R}}}_{\textnormal{cq}}=\widehat{{\mathcal{R}}}_{n+j}(y)$ defined in (8) for all $y\in{\mathcal{Y}}$.